The length (in cm) of 12 fishes caught from a lake is given as follows: 9.8, 9.1, 9.1, 11.3, 10.7, 10.2, 10.1, 9.7, 9.9, 9.5, 10.4, 10.5 Assume the population follows a normal distribution with variance of 4 cm. Calculate the 99% confidence interval to estimate the true mean length for all fishes in that lake


1
Expert's answer
2021-11-18T17:35:49-0500

"z=\frac{\overline{x}-\mu}{\sigma/\sqrt n}"

"\overline{x}=\sum x_i/n=10"

critical values for 99% confidence interval:

"z=\pm 2.576"


"-2.576<\frac{10-\mu}{2/\sqrt{12}}<2.576"


"-1.49<10-\mu<1.49"


"8.51<\mu<11.49"